# Efficient Paths for Local Counterdiabatic Driving

### v1

@@ Appendix D Explanation of Different Regimes of Fidelity Improvement
The exact matrix elements of the AGP in the energy eigenbasis satisfy
@@
We define $f(\omega)=-\sum_{k=1}^{\ell}\alpha_{k}\omega^{2k-1}$, where we have that
@@
Thus, finding the approximate AGP can be thought of as finding the best approximation to $1/\omega$ by a fixed number $(\ell)$ of odd polynomials. This fitting procedure is nicely illustrated in Figure 1 of Ref. Claeys et al., 2019. While this might seem at first glance to be a system-independent problem, it depends strongly on the range of excitation frequencies $\omega_{mn}$ which are present in the system.
@@
Figure 9: The variational optimization in terms of fitting excitation frequencies, for the short-range model with $N=12$. The top figure is the naive path, and the bottom is the augmented path. As before, this is just one snapshot when using extra control Hamiltonian $H_{c}=YY$. In this instance, it makes the local CD driving more effective by shifting the excitation frequencies to larger values, making them much easier to fit.
This fitting procedure happens for every point in the protocol. As an illustration we analyze here the point corresponding to $\lambda=0.5$, where the short-range model has a phase transition in the thermodynamic limit. In Figure 8, we show the polynomial $f(\omega)$ obtained for the short-range model with $N=6$ using the variational procedure together with the exact result $1/\omega$ using $H_{c}=YY$. Crucially, we plot only the transition frequencies corresponding to the excitations from the ground state, which we denote by $\omega_{m0}$, which have nonzero matrix element $\langle m|\partial_{\lambda}H|0\rangle$. This significantly reduces the number of relevant frequencies compared to the Hilbert space size. The number of points is further reduced because there are many nearly degenerate transition frequencies. As a result the AGP is only required to suppress a small number of transitions, and hence a small number of points to fit to a polynomial.

### v2

@@ Appendix E Explanation of Different Regimes of Fidelity Improvement
Figure 13: A snapshot of the variational optimization in terms of fitting excitation frequencies, for the short-range model with $N=12$. The top figure is the naive path, and the bottom is the augmented path. As before, this is just one snapshot when using extra control Hamiltonian $H_{c}=YY$. In this instance, it makes the local CD driving more effective by shifting the excitation frequencies to larger values (increasing the gap), making them much easier to fit.
The exact matrix elements of the AGP in the instantaneous energy eigenbasis satisfy
@@
Defining $f(\omega)=-\sum_{k=1}^{\ell}\alpha_{k}\omega^{2k-1}$, and trying to equate the matrix elements of the exact and approximate AGP gives
@@
Therefore, if the matrix elements exactly coincide then we have that $f(\omega_{mn})=1/\omega_{mn}$ i.e. we have fit $\omega_{mn}$ by $f(\omega_{mn})$ for all $\omega_{mn}$ present in the system. In other words, finding the approximate AGP can be thought of as finding the best approximation to $1/\omega$ by a fixed number $(\ell)$ of odd polynomials. This fitting procedure is nicely illustrated in Figure 1 of Ref. Claeys et al., 2019. While this might seem at first glance to be a system-independent problem, it depends strongly on the range of excitation frequencies $\omega_{mn}$ which are present in the system.
@@
This fitting procedure happens for every point in the protocol. As an illustration we analyze here the point corresponding to $\lambda=0.5$, where the short-range model has a phase transition in the thermodynamic limit. In Figure 12, we show the polynomial $f(\omega)$ obtained for the short-range model with $N=6$ using the variational procedure together with the exact result $1/\omega$ using $H_{c}=YY$. Crucially, we plot only the transition frequencies corresponding to the excitations from the ground state, which we denote by $\omega_{m0}$, which have nonzero matrix element $\langle m|\partial_{\lambda}H|0\rangle$. This significantly reduces the number of relevant frequencies compared to the Hilbert space size. The number of points is further reduced because there are many nearly degenerate transition frequencies. As a result the AGP is only required to suppress a small number of independent frequencies, which can be fit by a lower order polynomial perfectly.